# Domain Adaptation-Enhanced Searchlight: Enabling brain decoding from visual perception to mental imagery

### v1

@@ Abstract
In cognitive neuroscience and brain-computer interface research, accurately predicting imagined stimuli is crucial. This study investigates the effectiveness of Domain Adaptation (DA) in enhancing imagery prediction using primarily visual data from fMRI scans of 18 subjects. Initially, we train a baseline model on visual stimuli to predict imagined stimuli, utilizing data from 14 brain regions. We then develop several models to improve imagery prediction, comparing different DA methods. Our results demonstrate that DA significantly enhances imagery prediction, especially with the Regular Transfer approach. We then conduct a DA-enhanced searchlight analysis using Regular Transfer, followed by permutation-based statistical tests to identify brain regions where imagery decoding is consistently above chance across subjects. Our DA-enhanced searchlight predicts imagery contents in a highly distributed set of brain regions, including the visual cortex and the frontoparietal cortex, thereby outperforming standard cross-domain classification methods. The complete code and data for this paper have been made openly available for the use of the scientific community.
@@
Domain Adaptation , Brain Decoding , fMRI
journal: Neurocomputing
[…]
[inst1]organization=Intelligent Systems Group at the University of the Basque Country (UPV/EHU),city=San Sebastián, country=Spain
[…]
@@ 1 Introduction
Mapping brain activation patterns to their corresponding cognitive states is known as brain decoding. Methods for brain decoding are instrumental for understanding the brain processes involved in cognition [1] and therefore, an increasing number of works address the question of solving brain decoding tasks with machine learning algorithms [2, 3, 4].
@@
In the present study we hypothesize that failures to reveal similar neural representations may be due to the classification pipeline not accounting for the different distributions of the data across perception and imagery domains. Consequently, we claim that Domain Adaptation (DA) [11, 12] methods are an adequate approach to alleviate the hypothesized distribution shift and enhance the transferability of classifiers trained in visual perception to the imagery testing data.
Thus, we investigate the impact of different DA approaches in a particular fMRI cross-domain experiment where subjects were asked to solve perceptual and imagery tasks. Cross-domain scenarios involving imagery are the basis for Brain-Computer Interfaces (BCIs) [13] and have been addressed before [6, 14, 7, 15] but, to the best of our knowledge, the literature is scarce when it comes to applying DA techniques to such problems using fMRI data. Furthermore, rather than restricting our analysis to the visual cortex, we explore the whole brain using one local neighbourhood classifier per voxel (a searchlight procedure, as introduced by [16]). In this procedure, the brain is partitioned into thousands of overlapping spheres, and the voxels from each sphere are assigned to a classifier. The searchlight method combats the curse of dimensionality by limiting the radius of the sphere, and offers an anatomically interpretable output.
In this paper, we integrate a DA technique into the searchlight method to unveil common neural representations for perception and imagery across the brain.
In summary, the contributions made in this paper are:
@@
    First extensive evaluation and comparison of well-established DA techniques on cross-domain classification problems defined on fMRI data, considering several brain Regions Of Interest (ROIs).
@@
    Introducing the DA-enhanced searchlight approach using the best-performing DA technique on the whole brain, to gain more understanding on the nature and location of the domain-shift.
@@ 2 Methodological background
We define a domain $D$ as a combination of input and output spaces and an associated probability distribution: $(X,Y,p_{D})$. In the simplest case, there is a source domain $S$ with enough data to build a reliable learner, and we aim to generalize to a different target domain $T$ for which there is limited knowledge. In this paper, we consider the situation in which both domains share input and output spaces, meaning that they have the same dimensionality and live in the same space, but data come from different probability distributions $p_{S}$ and $p_{T}$
From the ML point of view, the problem we are interested in is risk minimization in the target domain, given a set of data $(x_{i},y_{i})\sim p_{S}(x,y)$ from the source distribution (following notation from [17]). Although the formulation is valid for a general output space $Y$, we will restrict ourselves to binary classification. In this setting, we would like to find the best learner $h^{\star}$, defined as the one with minimum expected loss $l$ under the target distribution, $h^{\star}=arg\underset{h\in\mathcal{H}}{min}R_{T}(h)$, where $h$ is any learner included in a given hypothesis space $\mathcal{H}$. The expression for the risk includes a term $p_{T}(x_{i},y_{i})$, that represents the probability of the source samples under the target distribution (Eq. 1, left hand side).
Introducing into the integral the probability of the observations coming from their actual source distribution $p_{S}$, we arrive to an expectation under $p_{S}$ of the loss multiplied by the ratio of target to source distributions (Eq. 1 right hand side).
@@
DA requires some previous knowledge of the target domain, be it in the form of some labeled (supervised DA, also known as few-shot learning) or unlabeled (unsupervised DA) instances from the target distribution. The amount of target domain knowledge incorporated into a DA technique is usually too scarce to build a model that relies in that information alone. In fact, for most DA applications, acquiring (or labeling) data from the target domain is difficult and/or expensive, but there is enough source domain data to build and validate a model with an adequate performance of unseen samples from that same source domain.
@@
1. Instance-based 1. Measure the difference between each source domain and target domain instances. 2. Adjust the domain instance contributions during training. 1. Real unequal eigenvalues of same sign 1. Balanced Weighting [20] 2. TrAdaBoost [21] 3. IWN [22] 1. Real unequal eigenvalues of same sign 1. KMM [23] 2. ULSIF [24] 3. RULSIF [25]
2. Feature-based 1. Look for a space of common features with respect to the task on source and target domain. 2. Build a representation under which $p_{S}$ and $p_{T}$ are similar. 2. Feature-based 1. FA and PRED [26] 2. Real unequal eigenvalues of opposite sign 1. SA [27]
3. Parameter-based 1. Adapt the parameters of a pre-trained source-only model to build a suitable model with the same structure on the target domain. 3. Equal eigenvalues Subtype 1: Two Independent vectors 1. Regular Transfer [28]
@@ 3 Related Work
In this section we review a number of works related to our research. We cover: 1) ML techniques for investigating commonalities in brain mechanisms for visual and imagery tasks; 2) DA methods applied to other neuroscience scenarios.
@@ 3.1 ML techniques for investigating commonalities in brain mechanisms for visual and imagery tasks
MVPA considers patterns of activity across multiple brain regions simultaneously, as opposed to univariate analysis which looks at each voxel independently. It is a widespread method for testing cognitive theories [29].
When it comes to the relatedness of perception and imagery, all published work in the decade of 2010 involves attempts of pure transfer using MVPA [6, 14, 30, 7]. These pioneering works evaluate a source-only classifier in the data from the target domain without implementing DA techniques. As we have already discussed, pure transfer is not guaranteed to work, due to a distribution shift between the brain patterns induced by different experimental conditions. Indeed, all of the aforementioned papers suffer from low generalization, showing a notable decrease between classifier performance in perception data and the attempted classification of imagery data.
@@
From 2023 onwards, deep learning methods have been employed to reconstruct mental images from brain activity patterns [34, 35], study visual object representation with a high spatiotemporal resolution [36] or learn robust representations from EEG recordings [37]. In [38], authors conduct a representational similarity analysis comparing the parameters of a CNN trained to recognize digits and a fMRI data-set containing brain patterns linked to digit recognition.
@@ 3.2 Other applications of DA in neuroscience
DA has been used in other neuroscientific settings, mostly in cross-device experiments for medical diagnosis. For instance, in [41], authors propose a low-rank DA method for autism diagnosis, aiming to determine a common lower-dimensional representation for data from multiple imaging devices. A very different approach involves the field of optimal transport, in which DA has been used to align the distributions of fMRI images from multiple devices with the goal of brain dementia identification [42]. Both of these methods fit into the category of feature-based DA, because they entail the projection of source and target domains to a new feature space where the domain shift has been minimized.
The use of instance-based approaches has also been reported in the literature. For example, in [43], the authors propose a re-weighting scheme for the diagnosis of Alzheimer’s disease. In the field of brain tumor segmentation, instance-based DA has been employed to correct sample selection bias amongst MRI image patches [44]. In [45], the authors compare the performance of four DA techniques in a cross-device MRI image segmentation task using support vector machines. Specifically, they review three instance-based approaches (including TrAdaboost [21]) and a parameter-based technique, similar to Regular Transfer [28].
@@ 4.1 Experimental setup
We quantified the contribution of 11 different DA algorithms to a binary cognitive state classification task using 18 data-sets, each containing around 9000-14000 anatomically selected voxels from approximately 500-600 fMRI images of a specific subject gathered in independent realizations of the same experiment, described in detail in [40]. The voxels come from 14 ROIs, the exact dimensionality of which is subject-dependent (Suppplementary Table 3). The ROIs were extracted using FreeSurfer [51], and their names are listed in Supplementary Table 2.
The experiment from which the data-sets were collected consisted on two phases: (1) perception and (2) imagery. In the perception step, each participant was presented with living (Dog) and non-living (Scissor) images from the THINGS [52] database. In the imagery stage, they received auditory cues instructing them to vividly imagine an object from one of those two categories for several seconds. The goal was to classify imagery instances into living or non-living, leveraging a classifier trained mostly on perception instances.
@@ 4.2 Machine learning task and validation framework
For each subject, we compared the performance of 11 DA approaches on the union of all ROIs using the scikit-learn compatible implementations in the ADAPT [53] library. B contains a brief description of each method.
As mentioned above, in the first place we evaluated the contribution of each DA technique to our task in the different subjects using the voxels from all 14 ROIs. Then, based on the comparison results, we also evaluated the performance of the best technique beyond the 14 ROIS, considering all the voxels from the whole brain. This whole-brain analysis was conducted using the searchlight [16] procedure.
@@ 4.3.1 Initial comparison of DA techniques
The evaluation criteria on which our comparisons are based are detailed hereafter. First, note that the methodology was constructed in such a way that the training of $h$, $h_{NAIVE}$ and $\{h_{DA}\}_{DA=1}^{11}$ were performed on exactly the same data, and the same is true for the evaluation phase of all estimators. During the 100 repetitions of the process, the balanced accuracies of each of the 13 estimators were measured on the test set of the target domain, obtaining the 13 corresponding performance vectors of size 100. Those 13 balanced accuracy vectors of $h$, $h_{NAIVE}$ and $\{h_{DA}\}_{DA=1}^{11}$ can be thought of as paired observations.
@@
Hence, our results are in the form of 10 balanced accuracy tables per subject, one corresponding to each value of $N_{t}$. Each table has 13 columns (one per ML approach) and 100 rows (one per data partition).
@@
We also concatenated all of the results (i.e. each of the 180 tables pertaining to each subject and value of $N_{t}$) to obtain a global comparison between all methods.
@@ 4.3.2 Evaluation of the searchlight analysis
For the searchlight analysis, we used the same data partitioning methodology as for the ROI-based analysis. In this case, each sample contains a whole-brain three-dimensional image with around 40,000 voxels (the exact number of which is subject-dependent - see Supplementary Table 3). Each voxel is a cube with a 3mm side, and we defined the radius of the searchlight spheres to be 12mm.
@@
    $h_{RTLC}$: Incorporating the same 100 samples from the target domain into $h$ via the RTLC algorithm.
@@
Then, we substracted 0.5 (the chance level balanced accuracy) from all the maps for further statistical testing.
@@
The statistical validation of our results employed the randomise technique [57] as implemented in FSL. This permutation-based strategy requires only minimal assumptions for validity (in our use case, exchangeability of the subjects and a symmetric distribution of the centered maps at a voxel level) and accounts for multiple comparisons when testing voxel by voxel [58]. Using this approach, we conducted a one sample t-test with variance smoothing ($\sigma=6$mm) and 10000 permutations. Threshold-Free Cluster Enhancement [59] was used, and the the family-wise error rate was controlled. This procedure yields a three-dimensional brain map of p-values, where each voxel indicates the probability that the balanced accuracy measured at said voxel across different subjects is statistically greater than chance. We chose a significance threshold of $\alpha=0.05$.
@@
5 Code and data availability
The code for this paper is available at the github repository AlexOlza/DA-enhanced-searchlight. The data is available at the Open Science Framework (OSF) platform at https://osf.io/audmx/.

### v3

@@ Abstract
In cognitive neuroscience and brain-computer interface research, accurately predicting imagined stimuli is crucial. This study investigates the effectiveness of Domain Adaptation (DA) in enhancing imagery prediction using primarily visual data from fMRI scans of 18 subjects. Initially, we train a baseline model on visual stimuli to predict imagined stimuli, utilizing data from 14 brain regions. We then develop several models to improve imagery prediction, comparing different DA methods. Our results demonstrate that DA significantly enhances imagery prediction in binary classification on our dataset, as well as in multiclass classification on a publicly available dataset. We then conduct a DA-enhanced searchlight analysis, followed by permutation-based statistical tests to identify brain regions where imagery decoding is consistently above chance across subjects. Our DA-enhanced searchlight predicts imagery contents in a highly distributed set of brain regions, including the visual cortex and the frontoparietal cortex, thereby outperforming standard cross-domain classification methods. The complete code and data for this paper have been made openly available for the use of the scientific community.
@@
Domain Adaptation, Brain Decoding, fMRI, Searchlight
@@ 1 Introduction
The mapping of brain activation patterns to their corresponding cognitive states is known as brain decoding. Methods for brain decoding are instrumental in understanding the brain processes involved in cognition [1] and therefore an increasing number of works address the question of solving brain decoding tasks with machine learning algorithms [2, 3, 4]. While “brain decoding” can encompass various tasks, including visual stimulus reconstruction, this study specifically focuses on the classification of neural patterns associated with visual perception and mental imagery.
@@
In the present study, we hypothesize that failures to reveal similar neural representations may be due to the classification pipeline not accounting for the different distributions of the data across perception and imagery domains. Consequently, we claim that Domain Adaptation (DA) methods [11, 12] are an adequate approach to alleviate the hypothesized distribution shift and enhance the transferability of classifiers trained in visual perception to imagery testing data.
Thus, we investigate the impact of different DA approaches in a particular fMRI cross-domain experiment where subjects were asked to solve perceptual and imagery tasks. Cross-domain classification tasks involving imagery are the basis for Brain-Computer Interfaces (BCIs) [13] and have been addressed before [6, 14, 7, 15] but, to our knowledge, the literature is scarce when it comes to applying DA techniques to such problems using fMRI data. Furthermore, rather than restricting our analysis to the visual cortex, we explore the whole brain using one local neighbourhood classifier per voxel (a searchlight procedure, as introduced by [16]). In this procedure, the brain is partitioned into thousands of overlapping spheres and the voxels of each sphere are assigned to a classifier. The searchlight method combats the curse of dimensionality by limiting the radius of the sphere, and offers an anatomically interpretable output.
In this paper, we integrate a DA technique into the searchlight method to unveil common neural representations for perception and imagery throughout the brain.
In summary, the contributions made in this paper are as follows.
@@
    First extensive evaluation and comparison of well-established DA techniques on cross-domain classification problems defined on fMRI data, considering several brain Regions Of Interest (ROIs) and two independent datasets.
@@
    Introducing the DA-enhanced searchlight approach using two different DA techniques on the whole brain, to gain more understanding on the nature and location of the domain-shift.
@@ 2 Methodological background
We define a domain $D$ as a combination of input and output spaces and an associated probability distribution: $(X,Y,p_{D})$. In the simplest case, there is a source domain $S$ with enough data to build a reliable learner, and our aim is to generalize to a different target domain $T$ for which there is limited knowledge. In this paper, we consider the situation in which both domains share input and output spaces, meaning that they have the same dimensionality and live in the same space, but data come from different probability distributions $p_{S}$ and $p_{T}$
From the ML point of view, the problem we are interested in is risk minimization in the target domain, given a set of data $(x_{i},y_{i})\sim p_{S}(x,y)$ from the source distribution (following notation from [17]). Although the formulation is valid for a general output space $Y$, we will restrict ourselves to binary classification. In this setting, we would like to find the best learner $h^{\star}$, defined as the one with minimum expected loss $l$ under the target distribution, $h^{\star}=arg\underset{h\in\mathcal{H}}{min}R_{T}(h)$, where $h$ is any learner included in a given hypothesis space $\mathcal{H}$. The expression for the risk includes a term $p_{T}(x_{i},y_{i})$, that represents the probability of the source samples under the target distribution (Eq. 1, left-hand side).
Introducing into the integral the probability of the observations coming from their actual source distribution $p_{S}$, we arrive to an expectation under $p_{S}$ of the loss multiplied by the ratio of target to source distributions (Eq. 1 right-hand side).
@@
DA requires some previous knowledge of the target domain, be it in the form of some labelled (supervised DA, also known as few-shot learning) or unlabelled (unsupervised DA) instances from the target distribution. The amount of target domain knowledge incorporated into a DA technique is usually too scarce to build a model that relies on that information alone. In fact, for most DA applications, acquiring (or labelling) data from the target domain is difficult and/or expensive, but there is enough source domain data to build and validate a model with an adequate performance of unseen samples from that same source domain.
@@
1. Instance-based Measure the difference between each source domain and target domain instances. Adjust the domain instance contributions during training. Balanced Weighting [20] TrAdaBoost [21] IWN [22] KMM [23] ULSIF [24] RULSIF [25]
2. Feature-based Look for a space of common features with respect to the task on source and target domain. Build a representation under which $p_{S}$ and $p_{T}$ are similar. 2. Feature-based FA and PRED [26] Fine-tuning [27] SA [28] MCD [29] DANN [30] Deep CORAL [31]
3. Parameter-based Adapt the parameters of a pre-trained source-only model to build a suitable model with the same structure on the target domain. Regular Transfer [32]
@@ 3 Related Work
In this section, we review a number of works related to our research. We cover: 1) ML techniques for investigating commonalities in brain mechanisms for visual and imagery tasks; 2) DA methods applied to other neuroscience scenarios.
@@ 3.1 ML techniques for investigating commonalities in brain mechanisms for visual and imagery tasks
MVPA considers patterns of activity across multiple brain regions simultaneously, as opposed to univariate analysis, which looks at each voxel independently. It is a widespread method for testing cognitive theories [33].
When it comes to the relatedness of perception and imagery, all published work in the decade of 2010 involves attempts of pure transfer using MVPA [6, 14, 34, 7]. These pioneering works evaluate a source-only classifier in the data from the target domain without implementing DA techniques. As we have already discussed, pure transfer is not guaranteed to work, due to a distribution shift between the brain patterns induced by different experimental conditions. Indeed, all the aforementioned papers suffer from low generalization, showing a notable decrease between classifier performance in perception data and the attempted classification of imagery data.
@@
From 2023 onwards, deep learning methods have been employed to reconstruct mental images from brain activity patterns [38, 39], study visual object representation with a high spatiotemporal resolution [40] or learn robust representations from EEG recordings [41]. In [42], the authors conduct a representational similarity analysis comparing the parameters of a CNN trained to recognize digits and a fMRI data-set containing brain patterns linked to digit recognition.
@@ 3.2 Other applications of DA in neuroscience
DA has been used in other neuroscientific settings, mostly in cross-device experiments for medical diagnosis. For instance, in [45], the authors propose a low-rank DA method for autism diagnosis, aiming to determine a common lower-dimensional representation for data from multiple imaging devices. A very different approach involves the field of optimal transport, in which DA has been used to align the distributions of fMRI images from multiple devices with the goal of brain dementia identification [46]. Both of these methods fit into the category of feature-based DA, because they entail the projection of source and target domains to a new feature space where the domain shift has been minimized.
The use of instance-based approaches has also been reported in the literature. For example, in [47], the authors propose a re-weighting scheme for the diagnosis of Alzheimer’s disease. In the field of brain tumour segmentation, instance-based DA has been employed to correct sample selection bias amongst MRI image patches [48]. In [49], the authors compare the performance of four DA techniques in a cross-device MRI image segmentation task using support vector machines. Specifically, they review three instance-based approaches (including TrAdaboost [21]) and a parameter-based technique, similar to Regular Transfer [32].
@@
The application of ML models to diverse human subjects beyond the training population is an important extension of DA in neuroscience. Cross-subject scenarios introduce additional complexity due to anatomical and functional variability between individual brains. Specifically, anatomical differences mean that each brain may have a distinct number of voxels, representing different feature spaces. Therefore, constructing a common feature space is essential before addressing the ML task. Traditional approaches for anatomical alignment across subjects rely on geometrical transformations, such as Tailarach coordinates, MNI space or hyperalignment techniques [55]. However, these methods have notable limitations, as they focus solely on anatomical alignment and do not account for functional variability in brain activity.
Recent studies have framed cross-subject learning as a multi-source DA problem, where each subject constitutes a separate source domain. Most advancements have targeted electroencephalogram (EEG) data, particularly for emotion recognition tasks [56, 57]. For example, in [56], the authors propose a two-level DA neural network (TDANN) preceded by a feature generator that maps each subject’s data into a unified space. Similarly, in [57], data from all subjects are first mapped into a shared feature space. Subsequent feature extractors are then used to capture subject-specific representations, and these representations’ joint distributions are aligned in a pairwise manner. However, advanced cross-subject DA methods for fMRI data remain underdeveloped. The high spatial resolution of fMRI data and the fact that each individual brain is different poses significant challenges to perform cross-subject DA. Additional work in this area would benefit from novel developments in hyperalignment algorithms [55] to use DA in a common high-dimensional voxel space across subjects.
@@ 4 Description of the cross-domain analysis
Figure 1 summarizes the experimental pipeline detailed in the following subsections.
Figure 1: Diagram of the experimental pipeline. Data acquisition: In the perception phase, fMRI data from 18 healthy subjects was acquired while they were presented with pictures of Living (Dog) and Non-living (Scissor) items. In the imagery phase, fMRI scans were recorded while the participants vividly imagined items from each category. DA comparison in the ROIs: The voxels from 14 anatomical Regions of Interest (ROIs) were used to compare 15 DA methods against a Logistic Regression (LR) baseline trained solely on perception and an alternative LR trained on perception and imagery. All the methods were tested on imagery. The comparison was repeated on a publicly available multiclass dataset. DA-enhanced searchlight: A local DA classifier was trained on the vicinity of each voxel in the brain, using data from perception and imagery, to produce a three-dimensional map of classifier accuracy. Created in BioRender. O, A. (2025) https://BioRender.com/u30c785
@@ 4.1 Experimental setup
We quantified the contribution of 15 different DA algorithms to a binary cognitive state classification task using 18 data-sets, each containing around 9000-14000 anatomically selected voxels from approximately 500-600 fMRI images of a specific subject gathered in independent realizations of the same experiment, described in detail in [44]. The voxels come from 14 ROIs, the exact dimensionality of which is subject-dependent (Supplementary Table 3). The ROIs were extracted using FreeSurfer [58], and their names are listed in Supplementary Table 2.
The experiment from which the data-sets were collected consisted of two phases: (1) perception and (2) imagery. In the perception step, each participant was presented with living (Dog) and non-living (Scissor) images from the THINGS [59] database. In the imagery stage, they received auditory cues instructing them to vividly imagine an object from one of those two categories for several seconds. The goal was to classify imagery instances into living or non-living, leveraging a classifier trained mostly on perception instances.
@@
Data preprocessing is thoroughly described in [44]. The first experimental volume, obtained after the heat-up phase, served as a functional reference. The rest of the volumes of each participant were co-registered to their corresponding reference volume. To minimize drift, the fMRI time-series was linearly detrended for each run. To capture peak neural activity in response to the presented stimuli, fMRI volumes from 5 to 18 seconds following the stimuli onset were retained (7 volumes per trial). The chosen samples were then independently stacked for region of interest (ROI) delineation. Next, the stacked perceptual volumes were z-scored by subtracting each voxel’s mean value and dividing by its standard deviation. Standardization of the imagery run scans was performed using the mean and standard deviation of the perceptual voxels to maintain a consistent and comparable scale across both conditions, facilitating analysis of the generalizability of information from perception to mental imagery.
We also performed the same analysis in a multiclass setting using the publicly available dataset “Generic Object Decoding” [35], that contains fMRI scans from 5 subjects in both the visual perception and imagery domains. The dataset description and the obtained results are shown in B.
@@ 4.2 Machine learning task and validation framework
For each subject, we compared the performance of 15 DA approaches on the union of all ROIs using the implementations in the ADAPT [60] library. D contains a brief description of each method. Due to the size of the dataset (with around 10 000 features), the Deep Learning methods (FT, DANN, MCD and DCORAL) required a prior dimensionality reduction step. Thus, we performed Independent Component Analysis (ICA) retaining 100 components.
As mentioned above, in the first place, we evaluated the contribution of each DA technique to our task in the different subjects using the voxels from all 14 ROIs. Then, based on the comparison results, we also evaluated the performance of the best technique beyond the 14 ROIs, considering all the voxels from the whole brain. This whole-brain analysis was conducted using the searchlight [16] procedure. To evaluate the robustness of the neuroscientific implications of our findings, we repeated this procedure using an alternative DA method.
@@ 4.3 Evaluation metrics and statistical validation
In this section, we detail the validation strategies, both for the initial comparison of DA techniques at the ROI level and for the DA-enhanced searchlight.
@@ 4.3.1 Initial comparison of DA techniques
The evaluation criteria on which our comparisons are based are detailed hereafter. First, note that the methodology was constructed in such a way that the training of $h$, $h_{NAIVE}$ and $\{h_{DA}\}_{DA=1}^{15}$ were performed on exactly the same data, and the same is true for the evaluation phase of all estimators. During the 100 repetitions of the process, the balanced accuracies of each of the 17 estimators were measured on the test set of the target domain, obtaining the 17 corresponding performance vectors of size 100. Those 17 balanced accuracy vectors of $h$, $h_{NAIVE}$ and $\{h_{DA}\}_{DA=1}^{15}$ can be thought of as paired observations.
@@
Hence, our results are in the form of 10 balanced accuracy tables per subject, one corresponding to each value of $N_{t}$. Each table has 17 columns (one per ML approach) and 100 rows (one per data partition).
@@
We also concatenated all the results (i.e. each of the 180 tables pertaining to each subject and value of $N_{t}$) to obtain a global comparison between all methods.
@@ 4.3.2 Evaluation of the searchlight analysis
For the searchlight analysis, we used the same data partitioning methodology as for the ROI-based analysis. In this case, each sample contains a whole-brain three-dimensional image with around 40,000 voxels (the exact number of which is subject-dependent - see Supplementary Table 3). Each voxel is a cube with a 3mm side, and we defined the radius of the searchlight spheres to be 12mm. We used algorithms RTLC and BW (see C for the results of the latter).
@@
    $h_{DA}$: Incorporating the same 100 samples from the target domain into $h$ via the RTLC or BW algorithm.
@@
Then, we subtracted 0.5 (the chance level balanced accuracy) from all the maps for further statistical testing.
@@
The statistical validation of our results employed the randomise technique [64] as implemented in FSL. This permutation-based strategy requires only minimal assumptions for validity (in our use case, exchangeability of the subjects and a symmetric distribution of the centred maps at a voxel level) and accounts for multiple comparisons when testing voxel by voxel [65]. Using this approach, we conducted a one sample t-test with variance smoothing ($\sigma=6$mm) and 10000 permutations. Threshold-Free Cluster Enhancement [66] was used, and the the family-wise error rate was controlled. This procedure yields a three-dimensional brain map of p-values, where each voxel indicates the probability that the balanced accuracy measured at said voxel across different subjects is statistically greater than chance. We chose a significance threshold of $\alpha=0.05$.